# A Methodology for Incompleteness-Tolerant and Modular Gradual Semantics for Argumentative Statement Graphs

### v6

@@ Appendix A Additional Properties
Appendix B contains theoretical results assessing the four structured GS against these four properties (they are incompatible with abstract GS). In summary, $\sigma_{T_{p}}$ and $\sigma_{T_{m}}$ satisfy Properties 14-17, while $\sigma_{\wedge_{D}}$ and $\sigma_{\wedge_{Q}}$ violate them. While each of these four properties differentiates our introduced GS from the existing GS, we believe this is reasonable given the properties’ focus on CSTs. Further, it can be seen that Properties 7 to 10 represent alternatives which guarantee suitable behaviour under all states of completeness (Definition 9).

### v7

@@ Appendix A Additional Properties
Appendix B contains theoretical results assessing the four structured GS against these four properties (they are incompatible with abstract GS). In summary, $\sigma_{T_{p}}$ and $\sigma_{T_{m}}$ satisfy Properties 14-17, while $\sigma_{\wedge_{D}}$ and $\sigma_{\wedge_{Q}}$ violate them.
While each of these four properties differentiates our introduced GS from the existing GS, we believe this is reasonable given the properties’ focus on CSTs. Further, it can be seen that Properties 7 to 10 represent alternatives which guarantee suitable behaviour under all states of completeness (Definition 9).